Logan Products computes its predetermined overhead rate annually on the basis of machine-hours. At the beginning of the year, it estimated that its total manufacturing overhead would be $349,500 and machines would be run a total of 20,500 hours. Its actual total manufacturing overhead for the year was $337,400 and its actual total machine-hours was 20,000 hours.

Required:

Compute the company’s predetermined overhead rate for the year, calculate the total overhead applied, and determine the amount of under- or overapplied overhead in the year. (Round your predetermined overhead rate to 2 decimal places.)

Predetermined overhead rate _____ per MH
Answer #1

a) Predetermine overhead rate = 349500/20500 = 17.05 per MH

Applied overhead = 20000*17.05 = $341000

Actual overhead = 337400

Over applied overhead = 341000-337400 = $3600